The Pritzker Music Pavilion in downtown Chicago is a technologically sophisticated and uniquely designed performing arts venue that hosts live concerts attended by over half a million patrons a year. A group of local organizers, led by a prominent local businesswoman, would like to use the pavilion for a concert to benefit a non-profit, national network of investors and environmental organizations working with companies and investors to address sustainability challenges such as global climate change. If the pavilion management agrees to host the concert, the organizers will donate all profits to the non-profit (or absorb any losses).

Based on the following revenue and cost information, the organizers would like answers to several questions.

There are three sources of revenue for the concert:

  1. Tickets will be sold for $14.50 each.
  2. A large multinational corporation headquartered in Chicago will donate $2.50 per ticket sold.
  3. Each concert attendee is expected to spend an average of $19.00 for parking, food, and merchandise.

On the expense side, there are also three components:

  1. A popular national group has agreed to perform at the concert. Normally, the group demands a significant fixed fee to perform, but to reduce the risk for the organizers, the group has agreed to perform for $5.50 per ticket sold.
  2. The organizers will pay several companies to operate the parking, food, and merchandise concessions. They will pay $22,000 plus 13% of all parking, food, and merchandise revenue.
  3. The organizers will pay the pavilion $80,000 plus $8.00 per ticket sold to cover its operating expenses (production, maintenance, advertising, etc.).

REQUIRED [ROUND YOUR CM ANSWER TO THE NEAREST CENT; ROUND ALL OTHER ANSWERS TO THE NEAREST UNIT OR NEAREST DOLLAR.]

Part A (8 tries; 8 points)
1. What is the estimated contribution margin per ticket sold for the benefit concert?   

2. What are the estimated total fixed costs for the benefit concert?   

Tries 0/8



Part B (8 tries; 8 points)
1. What is the estimated profit from the benefit concert if 8,500 tickets are sold?   

2. How many tickets must be sold in order for concert profit to be $90,000?   

3. Assuming a tax rate of 36% on profits from the concert, what must dollar ticket sales be in order for after-tax concert profits to be $90,000?   

Tries 0/8



Part C (4 tries; 4 points)
1. Assume that the organizers can negotiate the fixed portion of the pavilion's operating expenses. If the organizers expect to sell 8,500 tickets, how much operating fixed costs can they afford to pay and still earn a profit of $90,000 (ignore taxes)?
Answer #1

Part (a)

1. Contribution margin per ticket

Particulars Amount
Revenue per ticket (refer working note 1) 36.00
Less: - Variable Cost per ticket (refer working note 2) 15.97
Contribution margin per ticket 20.03

2. Calulation of Fixed Cost

Particulars Amount
Payment to companies as operating fees 22,000.00
Operating Expenses 80,000.00
Fixed Cost 1,02,000.00

Part (b)

1. Calculation of Estimated Profit:

Particulars Amount
Contribution fro sale of 8,500 tickets \left ( 20.03\times 8500 \right ) 1,70,255.00
Less: - Fixed Cost 1,02,000.00
Estimated Profit 68,255.00

2.  Calulation of number of units to be sold

Number of Unit to be sold =\frac{Fixed Cost+Desired Profit}{Contribution per ticket}

=\frac{102000+90000}{20.03}=9585.62\approx 9586

3.  Calulation of number of units to be sold

Number of Unit to be sold =\frac{Fixed Cost+Desired Profit(after tax)}{Contribution per ticket}

  102000+ VDDDDD 0.64 20.03 102000+ 140625 20.03 12113.08 12114

Part (c)

Calulation of Fixed Cost if change in offer

Number of Unit to be sold=\frac{Fixed Cost+Desired Profit}{Contribution per ticket}

8500=\frac{Fixed Cost+90000}{20.03}

170255=Fixed Cost+90000

Fixed Cost=170255-90000=80255

Fixed Cost to be paid to pavillion =$80255-$22000=$58255

Working Note 1:

Calculation of Revenue per ticket:

Particlars Amount
Sale price per ticket 14.50
Donation per ticket 2.50
Revenue from spending 19.00
Revenue per ticket 36.00

Working Note 2:

Calculation of Variable cost per ticket:

Particlars Amount
Fess to national group to perform at concert 5.50
Payment to companies as operating fees (=0.13\times 19=2.47) 2.47
Operating Expenses 8.00
Variable Cost per ticket 15.97